OpenStudy (anonymous):

determine whether the series is converges or diverges k=3 sigma k tends to infinity lnk/k

OpenStudy (anonymous):

diverges for sure the denominator is a polynomial of degree 1, so even if the numerator was a constant this would not converge

OpenStudy (anonymous):

how...?

OpenStudy (anonymous):

if you like you can use the "comparison test" and compare it to \(\sum\frac{1}{k}\)

OpenStudy (anonymous):

plz explain now what u sayy...?

OpenStudy (anonymous):

hellow comporisn test..?

OpenStudy (anonymous):

how can i do comporisn can u explain ..

OpenStudy (anonymous):

the denominator in this case is \(k\) and the numerator is \(\ln(k)\) this is larger than the well known divergent harmonic series \(\sum\frac{1}{k}\) since \(\sum\frac{1}{K}\) diverges, and since \(\frac{\ln(k)}{k}>\frac{1}{k}\) we know \(\sum \frac{\ln(k)}{k}\) diverges as well

OpenStudy (anonymous):

why ln(k)/k >1/k